How do you set up this word problem using I=P*R*T?
Steve has $25,000 to invest and wishes to earn an overall annual rate of return of 8%. His financial advisor recommends that he invest some of the money in a 5-year CD paying 4% per annum and the rest in a corporate bond paying 9% per annum. How much should be placed in each investment in order for Steve to achieve his goal?

Answer by stanbon(75887)

How do you set up this word problem using I=P*R*T?
Steve has $25,000 to invest and wishes to earn an overall annual rate of return of 8%. His financial advisor recommends that he invest some of the money in a 5-year CD paying 4% per annum and the rest in a corporate bond paying 9% per annum. How much should be placed in each investment in order for Steve to achieve his goal?
-------
Interest from the CD will be P*(0.05)(5)
Interest from the bond will be p(0.09)5
Equation:
interest + interest = interest
0.05*5x + 0.09*5*(25000-x) = 0.08*5*25000
-----------------
Multiply thru by 100 to get:
5*5x + 9*5(25000-x) = 8*525000
----
25x + 45*25000- 45x = 40*25000
-20x = -5*25000
x = $6250 (amt. to invest in 5% CDs)
----
25000-x = $18750 (amt. to invest in 9% bonds
=================================================
Cheers,
Stan H.